OpenStudy (anonymous):

Jake is comparing the prices of two mattress cleaning companies. Company A charges $30 per mattress and an additional $13 as service charges. Company B charges $28 per mattress and an additional $15 as service charges. Part A: Write equations to represent Company A and Company B's total mattress cleaning charges for a certain number of mattresses. Define the variable used in the equations. Part B: Which company would charge less for cleaning 4 mattresses? Justify your answer. Part C: How much money is saved by using the services of Company B instead of Company A to clea

OpenStudy (anonymous):

for company A x=30n+13 where x=the total price for A n=the matters number

OpenStudy (anonymous):

for company B y=28n+15 where y=the total price of B

OpenStudy (anonymous):

wow your good at this :)

OpenStudy (anonymous):

Is that it?

OpenStudy (anonymous):

the equations are correct but don't forget to define your variables. I like to use variables that match with what they stand for. So I would probably use P for final price, and M for # of mattresses. (but you can definitely use x&y since this is a linear equation. The service fee is your y intercept, they price per mattress is your slope) Company A charges $13 service fee (no matter how many mattresses) and $30 per mattress. So: P=30m+13 Do the same thing with company B: P=28m+15 Part B) Which company charges less for 4 mattresses? (solve for p when m=4) P=30(4)+13 P=120+13 P=133 Company A charges $133 for 4 mattresses OR P=28(4)+15 P=112+15 P=127 Company B charges $127 for 4 mattresses. So they have the better price in this situation.

OpenStudy (anonymous):

For part C I would need to know the rest of the question to help you answer it, but with the information provided you should be able to plug the numbers you were given to get the answer :) but just hollar if you need any more help!